# Scalable Decentralized Algorithms for Online Personalized Mean Estimation

### v3

@@ 6 Numerical Experiments
We evaluate the performance of the proposed algorithms over the class $\mathcal{G}_{0}(N,r)$ of simple regular graph, which models a situation where an agent is connected to $r$ other agents selected uniformly at random and provides the tree-like structure needed for the B-ColME as discussed in Sec. 5. The agents can belong to one of two similarity classes, associated with Gaussian distributions $\prescript{}{1}{D}\sim\mathcal{N}(\prescript{}{1}{\mu}=0,\sigma^{2}=4)$ and $\prescript{}{2}{D}\sim\mathcal{N}(\prescript{}{2}{\mu},\sigma^{2}=4)$. At initialization, each node is assigned one of the two classes with equal probability. Unless otherwise mentioned, in the experiments $|\mathcal{A}|=N=10000$, $r=10$, $d=4$, $\varepsilon=0.1$, $\delta=0.1$, and $\beta_{\gamma}(n)$ as per Eq. (3). In Appendix G we provide additional experiments varying the system’s parameters.
@@
Figure 1 showcases the performance of B-ColME and C-ColME by illustrating two key metrics: the fraction of agents with wrong estimates (more than $\varepsilon$ away from the true mean), and the fraction of wrong edges still used (a wrong edge is an edge connecting two agents belonging to a different class). We compare our algorithms against two benchmarks. In the first scenario, each agent independently relies on its local estimate, denoted as $\bar{x}^{t}_{a,a}$. In the second scenario, every agent $a$ is endowed with an oracle. This oracle precisely identifies which neighbors are in the same class as the agent, signified by $\mathcal{C}_{a}^{t}=\mathcal{C}_{a},\forall t$. The figure reveals that B-ColME has a longer transient phase but then exhibits a slightly steeper convergence than C-ColME. Notably, there is no apparent improvement in B-ColME’s estimates until about 90% of the incorrect edges have been removed, whereas C-ColME’s estimates begin to improve as soon as the first edges are eliminated. This phenomenon can be explained as follows: In B-ColME, the estimates at agent $a$ are not influenced by the removal of some incorrect edges as long as its $d$-hop neighborhood $\mathcal{N}_{a}^{d}$ remains unchanged. For instance, a given node $a^{\prime}\notin\mathcal{C}_{a}$ is removed from $\mathcal{N}_{a}^{d}$ only when all paths of length at most $d$ between agent $a$ and agent $a^{\prime}$ are eliminated. In contrast, in C-ColME, agent $a^{\prime}$ contributes to the weighted estimate at agent $a$ with a weight equal to the sum over all paths between $a$ and $a^{\prime}$ of the consensus coefficients along the path. As paths are progressively removed, the negative impact of $a^{\prime}$ on agent $a$’s estimate is gradually reduced. However, once all incorrect edges are removed, B-ColME benefits from its estimates being computed solely on agents belonging to the same class, while in C-ColME, some time is required for the effect of past estimates to fade away.
Figure 2: Comparison of our algorithms and the two versions of ColME, on a $G_{0}(N=10000,r=10)$ over 10 realizations.
@@
We also compare the proposed algorithms with ColME and a simplified version (s-ColME) in which the optimistic distance $d^{t}_{\gamma}(a,a^{\prime})$ is recomputed only for the $r$ agents that are queried at time $t$, achieving an $\mathcal{O}(r)$ per-agent computational cost (note that the memory cost remains $\mathcal{O}(|\mathcal{A}|)$). As predicted by the theoretical analysis, B-ColME and C-ColME are faster than ColME, but at the cost of a higher asymptotic error because agent $a$ collaborates only with the smaller group of nodes in $\mathcal{CC}_{a}^{d}$ in the case of B-ColME, and $\mathcal{CC}_{a}^{d}$ in the case of C-ColME. ColME pays for this asymptotic improvement with a $\mathcal{O}(|\mathcal{A}|)$ space-time complexity per agent, which makes the method impractical for large-scale systems. We observe that s-ColME improves ColME’s complexity at the cost of a much slower discovery of same-class neighbors, which significantly affects the estimation process.
While we focused on the fundamental problem of online mean estimation, our approach could be adapted to decentralized federated learning, for example by replacing the optimistic distance $d^{t}_{\gamma}(a,a^{\prime})$ with some estimator of the distance of the two local distributions. To illustrate this possibility, we consider a consensus-based decentralized federated learning algorithm Koloskova et al. [2020], where pairwise distances are computed on the basis of the cosine similarity of agents’ updates, similarly to what done in ClusteredFL Ghosh et al. [2020]; Chen et al. (details in Appendix H).
Figure 3 shows the performance of this algorithm considering $|\mathcal{A}|=100$ agents organized over a complete graph, and whose data is drawn from one of two different distributions. In our decentralized FL over a dynamic graph (FL-DG), agents exclude neighbors over time that they deem to belong to a different class, and the agent’s model is built only over similar neighbors, achieving higher accuracy.
@@ 7 Conclusions
In this paper, we introduced B-ColME and C-ColME, two scalable and fully distributed algorithms designed for collaborative estimation of the local mean. We have comprehensively evaluated the performance of these algorithms through both theoretical analysis and experimental validation. Our algorithms can be adapted to cooperatively learning personalized models within the context of federated learning.
Appendix A Appendix - Proof of Theorem 3
In this Appendix we provide the proof of Theorem 3 and, as a side result, we derive Equation (7). For the sake of clarity, we repeat the statement of the theorems. The proof of the results in Section 3.1 can be found in Asadi et al. [2022].
@@ Appendix A Appendix - Proof of Theorem 3
For any $\delta\in(0,1)$, employing B-ColME or C-ColME  we have:
@@
  • At last, combining previous results we can easily obtain the claim.
@@
($\sigma$-sub-Gaussian $x_{t}$) We start from the theoretical guarantees on the parametrization (Eq. 3) of the confidence intervals from Maillard [2019] [Lemma 2.7]: Let $\{x_{t}\}_{t\in\mathbb{N}}$ a sequence of independent real-valued random variables, where for each $t$, $x_{t}$ has mean $\mu_{t}$ and is $\sigma_{t}$-sub-Gaussian. Then, for all $\gamma\in(0,1)$ it holds:
@@
Our sequence of random variables can i) either correspond to the samples $x_{t}$ each node $a\in\mathcal{A}$ generates at each discrete time instant $t$, which is i.i.d., with mean $\mu_{a}$ and $\sigma$-sub-Gaussian (by assumption), or ii) the truncated sequence up to $n_{a,a^{\prime}}$ the node learns by querying its neighbors, possessing the same properties. Indeed, recall that for each locally available estimate $\bar{x}_{a,a^{\prime}}$, each node keeps also the number of samples over which that estimate has been computed $n^{t}_{a,a^{\prime}}$. For ease of notation we will drop the subscripts and superscripts, which for the sake of the lemma are superfluous. Being $\mu_{n^{\prime}}=\mu$ and $\sigma_{n^{\prime}}=\sigma$ constant in our case, it is immediate to write (considering just the first inequality for compactness):
@@
Dividing by $n$ both sides we obtain the sample mean $\bar{x}=\frac{1}{n}\sum_{t}^{n}x_{t}$ (in place of the summation) and the parametrization of confidence interval $\beta_{\gamma}(n)$, as we have introduced in Section 3 (which we restate here for completeness):
@@
$\leq$ $\frac{1}{n^{4}}[n\kappa\sigma^{4}+3n(n-1)\sigma^{4}]$
@@
$\mathbb{E}[(x^{t}-\mu)(x^{\tau}-\mu)(x^{\theta}-\mu)(x^{\phi}-\mu)]=\mathbb{E}[(x^{t}-\mu)]\mathbb{E}[(x^{\tau}-\mu)](x^{\theta}-\mu)(x^{\phi}-\mu)]=0$
@@
Now by sub-additivity of probability we get:
$\mathbb{P}(\exists n\in\mathbb{N},|\bar{x}-\mu|>\beta(n))=\mathbb{P}(\cup_{n}\{|\bar{x}-\mu|>\beta(n)\})<\sum_{n}\frac{\kappa\sigma^{4}}{n^{3}(\beta_{\gamma}(n))^{4}}+\sum_{n}\frac{3\sigma^{4}}{n^{2}(\beta_{\gamma}(n))^{4}}.$
@@
Imposing that this is smaller than $2\gamma$, we can concluded that by selecting
$\beta_{\gamma}(n)=\left(2\frac{\kappa+3\sigma^{4}}{\gamma}\right)^{\frac{1}{4}}\left(\frac{1+\ln^{2}n}{n}\right)^{\frac{1}{4}},$
@@
When $x^{t}$ exhibits a larger number of finite moments we can refine our approach by employing Proposition (11) for a different (larger) choice of $i$. So doing we obtain a more favorable behavior for $\beta_{\gamma}(n)$. In particular we will get that
@@
At last we wish to emphasize that $\beta_{\gamma}(n)$ can not be properly defined when distribution $D_{a}$ exhibit less than four bounded polynomial moments. The application of Chebyshev inequality ($i=1$), indeed, would lead to a too the following weak upper bound:
@@
Applying a union bound with respect to the nodes $a^{\prime}\in\mathcal{CC}_{a}$ and neighbors $a^{\prime\prime}\in\mathcal{N}_{a^{\prime}}$, and using Lemma 10 ($\mathbb{P}(\exists n\in\mathbb{N},|\bar{x}-\mu|\leqslant\beta_{\gamma}(n))\geq 2\gamma$), we can immediately obtain a lower bound on the probability of the event $E$:
@@
By Lemma 13 we have that, conditionally over $E$, as soon as $t-1\geq n^{\star}_{a,a^{\prime}}$ we have $d^{t}_{\gamma}(a,a^{\prime})>0$ for all pairs of neighboring nodes $(a,a^{\prime})$ belonging to different similarity classes, while $d^{t}_{\gamma}(a,a^{\prime})\leqslant 0$ for all pairs of neighboring nodes $(a,a^{\prime})$ belonging to the same similarity class. Therefore, whenever $t-1\leq n_{\gamma}^{\star}(a)$ this holds for all the pairs in the connected component $\mathcal{CC}_{a}$, as $\Delta_{a}\geq\min_{a^{\prime}\in{\mathcal{CC}_{a}}\setminus\mathcal{C}_{a}}\Delta_{a,a^{\prime}}$.
Appendix B Appendix - Schematic Representation for B-ColME

### v4

@@ 6 Numerical Experiments
We evaluate the proposed algorithms on the class $\mathcal{G}_{0}(N,r)$ of simple regular graphs (see Sec. 4.3). In this setting, each agent connects to $r$ other agents chosen uniformly at random in $\mathcal{A}$. This setup also provides the tree-like local structure required for B-ColME. Agents belong to one of two classes, associated with Gaussian distributions $D_{1}\sim\mathcal{N}(\mu_{1}=~{}0,\sigma^{2}=4)$ and $D_{2}\sim\mathcal{N}(\mu_{2}=1,\sigma^{2}=4)$. Each node is assigned to one of the two classes with equal probability. Unless otherwise stated, in the experiments $|\mathcal{A}|=N=10000$, $r=10$, $d=4$, $\varepsilon=0.1$, $\delta=0.1$, and $\beta_{\gamma}(n)$ as in (3). In Appendix C and I, we provide additional experiments for the multidimensional case and varying the system’s parameters.
@@
Figure 1 showcases the performance of B-ColME and C-ColME using two key metrics: the fraction of agents with incorrect estimates ($\hat{\mu}_{a}^{t}$ more than $\varepsilon$ away from the true mean $\mu_{a}$), and the fraction of wrong links still in use (a wrong link connects agents from different classes). We compare our algorithms against two benchmarks. The first benchmark has each agent independently relying on its local estimate $\bar{x}^{t}_{a,a}$. In the second benchmark, an oracle provides each agent with precise knowledge of which neighbors belong to the same similarity class (i.e., $\mathcal{C}_{a}^{t}=\mathcal{C}_{a}\cap\mathcal{N}_{a},\forall a,t$). The figure reveals that B-ColME has a longer transient phase but then exhibits a slightly steeper convergence than C-ColME. Notably, B-ColME’s estimates show no apparent improvement until about 90% of the wrong links have been removed, whereas C-ColME’s estimates begin to improve as soon as the first edges are eliminated. This phenomenon can be explained as follows. In B-ColME, the estimates at agent $a$ are not influenced by the removal of some wrong links as long as its $d$-hop neighborhood $\mathcal{N}_{a}^{d}$ remains unchanged. For instance, a given node $a^{\prime}\notin\mathcal{C}_{a}$ is removed from $\mathcal{N}_{a}^{d}$ only when all paths of length at most $d$ between agent $a$ and agent $a^{\prime}$ are eliminated. In contrast, in C-ColME, agent $a^{\prime}$ contributes to the weighted estimate at agent $a$ with a weight equal to the sum over all paths between $a$ and $a^{\prime}$ of the product of the consensus weights along the path. As paths are progressively removed, the negative impact of $a^{\prime}$ on agent $a$’s estimate is gradually reduced. However, once all wrong links are removed, B-ColME benefits from its estimates being computed solely on agents belonging to the same class, while C-ColME requires some additional time for the effect of past estimates to fade away.
Figure 2: Comparison of our algorithms and two versions of ColME, over 10 realizations.
@@
We also compare the proposed algorithms with ColME and a simplified version (s-ColME) where the optimistic distance $d^{t}_{\gamma}(a,a^{\prime})$ is recomputed only for the $r$ agents queried at time $t$, achieving an $\mathcal{O}(r)$ per-agent computational cost (the memory cost remains $\mathcal{O}(|\mathcal{A}|)$). As predicted by the theoretical analysis, B-ColME and C-ColME are faster than ColME, but at the cost of a higher asymptotic error because agent $a$ collaborates only with the smaller group of nodes in $\mathcal{CC}_{a}^{d}$ for B-ColME, and $\mathcal{CC}_{a}$ for C-ColME. ColME pays for this asymptotic improvement with a $\mathcal{O}(|\mathcal{A}|)$ space-time complexity per agent, impractical for large-scale systems. Note that s-ColME improves ColME’s complexity at the cost of a much slower discovery of same-class neighbors.
While we focused on online mean estimation, our approach can be adapted to decentralized federated learning. To illustrate this possibility, we adapt the consensus-based decentralized federated learning algorithm, by letting agents progressively exclude neighbors they identify as belonging to a different class. The cosine dissimilarity of agents’ updates, the same metric used in ClusteredFL (Ghosh et al. 2020; Chen et al. 2021), replaces the optimistic distance $d^{t}_{\gamma}(a,a^{\prime})$ (details in Appendix J). Figure 3 shows the performance of our decentralized FL over a dynamic graph (FL-DG) with $|\mathcal{A}|=100$ agents initially organized over a complete graph. Two different distributions are obtained from MNIST (Deng 2012) by swapping/maintaining some labels and each client progressively receives new data samples from one of the two distributions. As the graph is progressively split in two clusters of clients belonging to the same class, each agent’s model benefits from cooperating only with similar clients and it achieves a higher accuracy.
Figure 3: Accuracy of a local model (Local), a decentralized FL over a static graph (FL-SG), and our approach over a dynamic graph (FL-DG). We also show the fraction of links between classes (wrong links) over time for FL-DG.
@@ 7 Conclusions
In this paper, we introduced B-ColME and C-ColME, two scalable and fully distributed algorithms for collaborative local mean estimation. We thoroughly evaluated their performance through both theoretical and empirical analyses. Additionally, we adapted our approach for personalized federated learning, applying it to the task of handwritten digit recognition using the MNIST dataset.
This work points to several future research directions. Here we have allowed agents only to sever existing connections, but not to establish new ones. Investigating scenarios where agents can rewire their connections to communicate with new agents outside their original neighborhood would be an interesting extension. Additionally, we assumed that agents are partitioned into similarity classes, with agents in the same class generating data with identical true mean. Extending our approach to accommodate more general scenarios, where each agent generates data with potentially different true mean, would be a valuable avenue for further exploration.
Acknowledgements
This research was supported in part by the European Network of Excellence dAIEDGE under Grant Agreement Nr. 101120726 and by the Groupe La Poste, sponsor of the Inria Foundation, in the framework of the FedMalin Inria Challenge. It was also funded by the Nokia-Inria challenge LearnNet, and by the French government National Research Agency (ANR) through the UCA JEDI (ANR-15-IDEX-0001), EUR DS4H (ANR-17-EURE-0004), and the 3IA Côte d’Azur Investments in the Future project (ANR-19-P3IA-0002).
Computational resources provided by hpc@polito,which is a project of Academic Computing within the Department of Control and Computer Engineering at the Politecnico di Torino (http://hpc.polito.it).
Appendix A Appendix A - ColME’s Guarantees
We provide the two main theoretical results from Asadi et al. (2022) [Theorem 1 and Theorem 2] which were proven under the assumption that the distributions $\{D_{a},a\in\mathcal{A}\}$ are sub-Gaussian. Recall that a distribution $D$ is sub-Gaussian with parameter $\sigma^{2}$ if $\forall\lambda\in\mathbb{R}$, $\log\mathbb{E}_{x\sim D}\exp(\lambda(x-\mu))\leqslant\frac{1}{2}\lambda^{2}\sigma^{2}$. The proofs of the Theorems presented in this Appendix can be found in (Asadi et al. 2022).
Select the amplitude of the confidence intervals $I_{a,a}$ and $I_{a,a^{\prime}}$, centered around the sample means, as:
Theorem 7.
(Asadi et al. 2022, Theorem 1) Assume distributions $\{D_{a},\forall a\in\mathcal{A}\}$ are sub-Gaussians with parameter $\sigma^{2}$. For any $\delta\in(0,1)$, and with $\gamma=\frac{\delta}{4|\mathcal{A}|}$, employing ColME, we have:
Turning our attention to the estimation error, it holds:
Theorem 8.
(Asadi et al. 2022, Theorem 2) Assume distributions $\{D_{a},\forall a\in\mathcal{A}\}$ are sub-Gaussians with parameter $\sigma^{2}$. For any $\delta\in(0,1)$, and with $\gamma=\frac{\delta}{4|\mathcal{A}|}$, employing ColME, we have:
Appendix B Appendix B - Proof of Theorem 1
In this Appendix we provide the proof of Theorem 1 and, as a side result, we derive Equation (5). For the sake of clarity, we repeat the statement of the theorems.
@@ Appendix B Appendix B - Proof of Theorem 1
Considering an agent $a$ picked arbitrarily in $\mathcal{A}$, for any $\delta\in(0,1)$, employing B-ColME or C-ColME  we have:
@@
  • At last, by combining previous results we can easily obtain the claim.
@@
($\sigma$-sub-Gaussian $x_{t}$) We start from the theoretical guarantees on the parametrization (Eq. 10) of the confidence intervals from (Maillard 2019) [Lemma 2.7]: Let $\{x_{t}\}_{t\in\mathbb{N}}$ a sequence of independent real-valued random variables, where for each $t$, $x_{t}$ has mean $\mu_{t}$ and is $\sigma_{t}$-sub-Gaussian. Then, for all $\gamma\in(0,1)$ it holds:
@@
Our sequence of random variables can i) either correspond to the samples $x_{t}$ each node $a\in\mathcal{A}$ generates at each discrete time instant $t$, which is i.i.d., with mean $\mu_{a}$ and $\sigma$-sub-Gaussian (by assumption), or ii) the truncated sequence up to $n_{a,a^{\prime}}$ the node learns by querying its neighbors, possessing the same properties. Indeed, recall that for each locally available estimate $\bar{x}_{a,a^{\prime}}$, each node keeps also the number of samples over which that estimate has been computed $n^{t}_{a,a^{\prime}}$. For ease of notation, we will drop the subscripts and superscripts, which for the sake of the lemma are superfluous. Being $\mu_{n^{\prime}}=\mu$ and $\sigma_{n^{\prime}}=\sigma$ constant in our case, it is immediate to write (considering just the first inequality for compactness):
@@
Dividing by $n$ both sides we obtain the sample mean $\bar{x}=\frac{1}{n}\sum_{t}^{n}x_{t}$ (in place of the summation) and the parametrization of confidence interval $\beta_{\gamma}(n)$, as we have introduced in Sec. 3 (which we restate here for completeness):
@@
$\leq$ $\frac{1}{n^{4}}[n\kappa\sigma^{4}+3n(n-1)\sigma^{4}]$ (18)
@@
$\mathbb{E}[(x^{t}-\mu)(x^{\tau}-\mu)(x^{\theta}-\mu)(x^{\phi}-\mu)]=\mathbb{E}[(x^{t}-\mu)]\mathbb{E}[(x^{\tau}-\mu)(x^{\theta}-\mu)(x^{\phi}-\mu)]=0$
@@
Now by sub-additivity of probability, we get:
$\mathbb{P}(\exists n\in\mathbb{N},|\bar{x}-\mu|>\beta_{\gamma}(n))=\mathbb{P}(\cup_{n}\{|\bar{x}-\mu|>\beta_{\gamma}(n)\})<\sum_{n}\frac{\kappa\sigma^{4}}{n^{3}(\beta_{\gamma}(n))^{4}}+\sum_{n}\frac{3\sigma^{4}}{n^{2}(\beta_{\gamma}(n))^{4}}.$
@@
Imposing that this is smaller than $2\gamma$, we can conclude that by selecting
$\beta_{\gamma}(n)=\left(2\frac{(\kappa+3)\sigma^{4}}{\gamma}\right)^{\frac{1}{4}}\left(\frac{1+\ln^{2}n}{n}\right)^{\frac{1}{4}},$
@@
When $x^{t}$ exhibits a larger number of finite moments we can refine our approach by employing Proposition (11) for a different (larger) choice of $i$. Doing so, we obtain a more favorable behavior for $\beta_{\gamma}(n)$. In particular, we will get that
@@
At last, we wish to emphasize that $\beta_{\gamma}(n)$ can not be properly defined when distribution $D_{a}$ exhibits less than four bounded polynomial moments. The application of Chebyshev inequality ($i=1$), indeed, would lead to a too the following weak upper bound:
@@
Applying a union bound with respect to the nodes $a^{\prime}\in\mathcal{CC}_{a}$ and neighbors $a^{\prime\prime}\in\mathcal{N}_{a^{\prime}}$, and using Lemma 10 ($\mathbb{P}(\exists n\in\mathbb{N},|\bar{x}-\mu|\geq\beta_{\gamma}(n))\leq 2\gamma$), we can immediately obtain a lower bound on the probability of the event $E$:
@@
By Lemma 13 we have that, conditionally over $E$, as soon as $t-1\geq n^{\star}_{a,a^{\prime}}$ we have $d^{t}_{\gamma}(a,a^{\prime})>0$ for all pairs of neighboring nodes $(a,a^{\prime})$ belonging to different similarity classes, while $d^{t}_{\gamma}(a,a^{\prime})\leqslant 0$ for all pairs of neighboring nodes $(a,a^{\prime})$ belonging to the same similarity class. Therefore, whenever $t-1\geq n_{\gamma}^{\star}(a)$ this holds for all the pairs in the connected component $\mathcal{CC}_{a}$, as $\Delta_{a}\geq\min_{a^{\prime}\in{\mathcal{CC}_{a}}\setminus\mathcal{C}_{a}}\Delta_{a,a^{\prime}}$. ∎
Appendix C Appendix C - Extension to the Multidimensional Case
In this Appendix, we show briefly how the previous approach can be generalized to the multidimensional case. We assume that at each instant $t$ every agent $a$ generates a new sample $\bm{x}_{a}^{t}\in\mathbb{R}^{K}$ drawn i.i.d. from distribution $D_{a}$ with expected value $\bm{\mu}_{a}=\mathbb{E}[\bm{x}_{a}^{t}]$. In particular, we show how the definition of confidence intervals $\beta_{\gamma}(n)$ can be extended to the multidimensional case. Let $\Delta_{a,a^{\prime}}:=||\bm{\mu}_{a}-\bm{\mu}_{a^{\prime}}||$.
C.1 Sub-Gaussian case
In this case, we assume distributions $D_{a}$ to be multidimensional sub-Gaussians. Recalling the definition of multidimensional sub-Gaussian random variables, we have:
Definition 1.
A random vector $\bm{x}\in\mathbb{R}^{K}$ is said to be sub-Gaussian with variance proxy $\sigma^{2}$ if it is centered and for any $\bm{u}\in\mathbb{R}^{K}$ such that $||u||=1$, the real random variable $\bm{u}^{T}\bm{x}$ is sub-Gaussian with variance proxy $\sigma^{2}$.
Therefore, we can choose an orthonormal basis $\bm{u}_{i}$ $\forall 1\leq i\leq K$ in $\mathbb{R}^{K}$ and consider the uni-dimensional projections of samples along directions induced by $\bm{u}_{i}$, $\forall i$. Since projected samples are uni-dimensional sub-Gaussian, the theory developed in the previous sections applies to every projection $i$. In particular, we can compute the optimistic distance along axes $i$ as:
$d_{i,\gamma/K}^{t}(a,a^{\prime}):=\left|\bm{u}_{i}^{T}(\bar{\bm{x}}_{a,a}^{t}-\bar{\bm{x}}_{a,a^{\prime}}^{t})\right|-\beta_{\gamma/K}(n_{a,a}^{t})-\beta_{\gamma/K}(n_{a,a^{\prime}}^{t})$
and decide that $a^{\prime}$ is maintained in $\mathcal{C}^{t}_{a}$, as long as $d_{i,\gamma/K}^{t}(a,a^{\prime})\leq 0$  $\forall\;1\leq i\leq d$. At last observe that $\Delta^{(i)}_{a,a^{\prime}}=\bm{u}_{i}^{T}(\bm{\mu}_{a}-\bm{\mu}_{a^{\prime}})$ satisfies:
$\max_{i}\Delta^{(i)}_{a,a^{\prime}}\geq\frac{1}{\sqrt{K}}\Delta_{a,a^{\prime}}$
as a result of simple geometrical arguments.
C.2 Fourth bounded moment
We start assuming that that $\mathbb{E}[||(\bm{x}_{a}^{t}-\bm{\mu}_{a})||^{4}]<\infty$ for any $a\in\mathcal{A}$. Let $\bar{\bm{x}}_{a}=\frac{1}{n}\sum_{t=1}^{n}\bm{x}^{t}$ and $Y_{a}:=||\bar{\bm{x}}_{a}-\bm{\mu}_{a}||\geq 0$, by applying Proposition 11 to $Y$ we have:
$\mathbb{P}(|Y_{a}|>\beta_{\gamma}(n))\leq\frac{\mathbb{E}[Y_{a}^{4}]}{\beta_{\gamma}(n)^{4}}$
Now defined with ${\mu}_{i,a}$ the expectation of the $i$-th component of samples $\bm{x}^{t}_{a}$, i.e:
$\mu_{i,a}:=\mathbb{E}[x^{t}_{i,a}]$
considering the Euclidean norm, we have:
$\mathbb{E}[Y_{a}^{4}]$ $=\mathbb{E}\left[\left(\sum_{i=1}^{K}(\bar{x}_{i,a}-{\mu}_{i,a})^{2}\right)^{2}\right]=\sum_{i=1}^{K}\sum_{h=1}^{K}\mathbb{E}[(\bar{{x}}_{i,a}-\mu_{i,a})^{2}({\bar{x}}_{h,a}-\mu_{h,a})^{2}]$
$\leq\sum_{i=1}^{K}\sum_{h=1}^{K}\sqrt{\mathbb{E}[(\bar{{x}}_{i,a}-\mu_{i,a})^{4}]\mathbb{E}[(\bar{{x}}_{h,a}-\mu_{h,a})^{4}]}\leq K^{2}\max_{i}\mathbb{E}[(\bar{{x}}_{i,a}-\mu_{i,a})^{4}]$
where the second to last inequality follows by the application of the Cauchy-Schwarz inequality.
Now proceeding exactly as in (18) we obtain that for every $i$
$\mathbb{E}[(\bar{{x}}_{i,a}-\mu_{i,a})^{4}]\leq\frac{1}{n^{4}}[n\kappa_{i}\sigma_{i}^{4}+3n(n-1)\sigma_{i}^{4}]$
Now by sub-additivity of probability we get, recalling that $Y_{a}=||\bar{\bm{x}}_{a}-\bm{\mu}_{a}||$:
$\mathbb{P}(\exists n\in\mathbb{N},Y_{a}>\beta_{\gamma}(n))=\mathbb{P}(\cup_{n}\{Y_{a}>\beta_{\gamma}(n)\})<K^{2}\max_{i}\left(\sum_{n}\frac{\kappa_{i}\sigma^{4}_{i}}{n^{3}(\beta_{\gamma}(n))^{4}}+\sum_{n}\frac{3\sigma_{i}^{4}}{n^{2}(\beta_{\gamma}(n))^{4}}\right).$
Then proceeding as in the proof of Theorem 1, (i.e. forcing the r.h.s to be less than $2\gamma$) we obtain the following expression for $\beta_{\gamma}(n)$:
$\beta_{\gamma}(n)=\max_{i}\sqrt{K}\left(2\frac{(\kappa_{i}+3)\sigma_{i}^{4}}{\gamma}\right)^{\frac{1}{4}}\left(\frac{1+\ln^{2}n}{n}\right)^{\frac{1}{4}}.$
At last, we have to generalize the definition of optimistic distance as follows:
$d_{\gamma}^{t}(a,a^{\prime}):=\left|\left|\bar{\bm{x}}_{a,a}^{t}-\bar{\bm{x}}_{a,a^{\prime}}^{t}\right|\right|-\beta_{\gamma}(n_{a,a}^{t})-\beta_{\gamma}(n_{a,a^{\prime}}^{t}).$
Then, proceeding exactly as in the previous section, we can extend Lemma 13 statement to the multidimensional case (we recall that by triangular inequality $\left|\left|\bar{\bm{x}}_{a,a}^{t}-\bar{\bm{x}}_{a,a^{\prime}}^{t}\right|\right|=\left|\left|(\bm{\mu}_{a}-\bm{\mu}_{a^{\prime}})+(\bar{\bm{x}}_{a,a}^{t}-\bm{\mu}_{a})-(\bar{\bm{x}}_{a,a^{\prime}}^{t}-\bm{\mu}_{a^{\prime}})\right|\right|\leq\Delta_{a,a^{\prime}}+\left|\left|\bar{\bm{x}}_{a,a}^{t}-\bm{\mu}_{a}\right|\right|+\left|\left|\bar{\bm{x}}_{a,a^{\prime}}^{t}-\bm{\mu}_{a^{\prime}}\right|\right|$, as well as by reverse triangular inequality $\left|\left|\bar{\bm{x}}_{a,a}^{t}-\bar{\bm{x}}_{a,a^{\prime}}^{t}\right|\right|\geq\Delta_{a,a^{\prime}}-\left|\left|\bar{\bm{x}}_{a,a}^{t}-\bm{\mu}_{a}\right|\right|-\left|\left|\bar{\bm{x}}_{a,a^{\prime}}^{t}-\bm{\mu}_{a^{\prime}}\right|\right|$).
C.3 Numerical Experiments
We experimentally assess the ‘dicovery’ phase, where nodes attempt to find peers in their neighborhood $\mathcal{N}_{a}$ that belong to the same similarity class (note that the guarantees we established for the estimate process, whether through the message-passing scheme or consensus, hold for each component).
We consider a setup similar to Section 6, with $N=10000$ agents, belonging to one of two classes: the first characterized by an all-zero vector mean $\bm{\mu_{1}}=\mathbf{0}$ and the second by $\bm{\mu_{2}}=\frac{1}{\sqrt{K}}\mathbf{1}$. Samples are drawn from a multivariate Gaussian distribution, where the common variance (the diagonal element in the covariance matrix) is $\sigma^{2}=4.0$. We consider the canonical basis , i.e., $\{\mathbf{e}_{i}\}_{i\in\{1,...,K\}}$  [^5] The vector $\mathbf{e}_{i}=(0,...,1,...,0)$ is the all-zero vector with a a $1$ in its i-th component., as the vectors for projecting our vectorial samples. The covariance matrix can be arbitrary, provided that the diagonal elements (the variances) are $\sigma^{2}$. We consider $\varepsilon=0.1$ and $\delta=0.1$, averaging over $10$ different realizations of the process.
Figure 4: Fraction of the wrong links (log scale) over time, as a function of the dimension $K$ of the mean values, $\bm{x}\in\mathbb{R}^{K}$.
As the number of dimensions $K$ increases, the time required to discover peers in the same class also increases. However, doubling the number of dimensions results in less than a doubling of the discovery time. This slowdown occurs because the bounds on each component need to be more stringent (note the factor $\frac{\gamma}{K}$ in Sec. C.1) to guarantee the same bound on the probability of incorrectly hindering a connection to a same-class peer. However, it’s also important to note that if any of the $K$ components of the distance $d_{\gamma/K}^{t}(a,a^{\prime})$ becomes positive, it is a condition sufficient to remove the incorrect link. This mechanism contributes to the sublinear increase in discovery time with respect to $K$.
Appendix D Appendix D - Schematic Representation for B-ColME